# Auditing Meta-Cognitive Hallucinations in Reasoning Large Language Models

### v2

@@ B.1 Dataset Overview
Our Controlled Hallucination Audit Dataset, the first to audit Long-CoT hallucinations in RLLMs, primarily comprises question and reasoning-answer generation. All data synthesis was conducted under strict human oversight to ensure annotation quality. The dataset is divided into four subsets: Type I (Seen but Unlearned), Type I Control (Correct Answer), Type II (Unseen or Erroneous), and Type II Control (Error Rejected). All RLLMs employed for question and reasoning–answer synthesis utilize DeepSeek-R1 [39]. We chose DeepSeek-R1 as the main tested model because its characteristics align well with the needs of our study. Specifically, the model’s relatively high hallucination rate provides a fertile ground for systematic error analysis, while its low inference cost enables large-scale experimentation. In addition, its strong recognition within the open-source community ensures reproducibility and broad relevance. These properties make DeepSeek-R1 an appropriate foundation for controlled analysis of reasoning errors in CoT-style outputs (see Appendix B.2 for comparative results on other models).The following presents the dataset’s construction principles and workflow.
@@ E.1 Evaluation Setting
Model Setting. For the Dataset Construction section, we used the DeepSeek-R1 API [39] and ChatGPT-4o API [2] to synthesize data and assist in the manual verification of samples. We evaluated performance based on DeepSeek’s officially released distilled model, DeepSeek-R1-Distill-Qwen-14B [39] for the Hallucination Detection section. Some detection methods required long-text understanding and processing, such as splitting full CoTs into individual claims, for which we employed DeepSeek-V3 API [39] and ChatGPT-4o API.
@@ E.2.1 Logit-Based Detection
Logit entropy measures the dispersion of probability mass across the most likely next tokens, making it effective at revealing the model’s uncertainty and exposing potential factual inconsistencies. The model’s output logits are first converted into token‐level perplexity, with higher perplexity indicating greater uncertainty or likelihood of hallucination. For detection, we compute top-k logit entropy by normalizing the entropy over the K most probable tokens, following Sriramanan et al. [52], we select an optimal threshold on the validation set, and then apply that threshold during testing to flag hallucinated outputs.
@@ E.2.2 Attention-Based Detection
The attention-kernel score is computed by taking the logarithm of each token’s self-attention weight (i.e., the diagonal entries of the attention matrix) and averaging across tokens. This score effectively measures the model’s degree of “self-focus” during reasoning: consistently high self-attention suggests the model is coherently building on its previous inference, whereas a weakened diagonal focus may indicate hallucination. Following the approach of [52], we calculate the attention-kernel score layer by layer and determine, on a validation set, the optimal threshold for each layer. We apply the threshold corresponding to the layer that achieved the best validation performance for testing.
@@ E.2.3 Hidden-State Based Detection
For each sample (question–answer pair), we obtain the hidden representations of each layer in teacher-forcing mode. For each layer’s activation matrix, we compute the centered covariance and perform singular value decomposition (SVD), then average the logarithm of the singular values to derive that layer’s hidden-state score. Higher scores indicate more complete, structured, and reliable internal representations; lower scores often correspond to degraded or incoherent features, potentially signaling hallucination. Following Sriramanan et al. [52], we determine the optimal threshold for each layer on a balanced validation set, and during testing, we apply the threshold from the best-performing layer to detect hallucinations in new samples.
@@ E.2.4 HalluciNot (HDM2 model)
The HDM2 model [48] detects hallucinations by verifying both context-specific facts and common‐knowledge statements against a fine‐grained taxonomy of LLM outputs. Given a user prompt, an optional context, and an LLM response, it first categorizes each sentence into one of four classes (context‐based, common‐knowledge, enterprise‐specific, or innocuous) using a lightweight classifier. For context‐based claims, it employs a retrieval‐augmented consistency check. For common‐knowledge claims, it probes a frozen backbone LLM’s internal representations to identify contradictions with widely accepted facts. It then produces a document‐level hallucination score and a group of token‐level scores. We took document‐level hallucination scores to evaluate hallucinations.
@@ E.2.5 Claim-Conditioned Perplexity (CCP)
The CCP method quantifies the model’s confidence in factual statements and can be used to detect latent hallucinations in generated text. First, the model’s output is segmented into atomic factual claims, each representing an independent fact unit. We retrieve each token in a claim’s probability distribution over the top-K candidates given the context. Using a natural language inference (NLI) model, we filter these candidates to retain only those that entail or contradict the original token, and compute a normalized ratio of their probabilities as the token’s confidence score. We then aggregate the token-level confidence scores within each claim to obtain an overall uncertainty score for that claim. We average the uncertainty scores across all claims to evaluate an entire text. Finally, we determine an optimal threshold on a validation set to decide whether the text contains hallucinations. We follow the approach described in [18] to perform the computations.
@@ E.2.6 SelfCheckGPT
SelfCheckGPT evaluates a model’s self-consistency under diverse random sampling to detect latent hallucinations. Given a user prompt, it first generates a deterministic “main” response at temperature 0, then produces $N$ stochastic samples at temperature 1. The main response is segmented into sentences, and each sentence is compared against all sampled outputs via multiple consistency checks, such as maximal BERTScore similarity inversion, NLI-based contradiction probability, n-gram model token scoring, and prompt-based “Does this hold?” queries. Each check yields a per-sentence uncertainty score, which are aggregated to form both sentence-level and response-level hallucination scores. A threshold tuned on a validation set is then applied to classify sentences or entire responses as hallucinated.We follow the procedure in [42] to perform these calculations.
@@ E.2.7 Semantic Entropy
The Semantic-Entropy score assesses semantic uncertainty by measuring the entropy of the answer distribution obtained through multiple samplings of adversarial prompts. First, the generated text is segmented into atomic claims, and several interrogative prompts are automatically constructed for each claim. Each prompt is then sampled multiple times to gather all answers along with their generation probabilities. Using a bidirectional entailment model, answers are clustered by semantic consistency, and the probabilities within each cluster are summed to form a cluster-level distribution. The information entropy of this distribution is computed as the semantic uncertainty metric for that prompt. We then average the entropies of all prompts associated with the same claim to derive the claim’s overall semantic entropy. Finally, an optimal entropy threshold is determined on the validation set and applied during testing to detect hallucinations. We follow the procedure in [19] to perform these calculations.

### v3

@@ B.1 Dataset Overview
Our Controlled Hallucination Audit Dataset, the first to audit Long-CoT hallucinations in RLLMs, primarily comprises question and reasoning-answer generation. All data synthesis was conducted under strict human oversight to ensure annotation quality. The dataset is divided into four subsets: Type I (Seen but Unlearned), Type I Control (Correct Answer), Type II (Unseen or Erroneous), and Type II Control (Error Rejected). All RLLMs employed for question and reasoning–answer synthesis utilize DeepSeek-R1 [liu2024deepseek]. We chose DeepSeek-R1 as the main tested model because its characteristics align well with the needs of our study. Specifically, the model’s relatively high hallucination rate provides a fertile ground for systematic error analysis, while its low inference cost enables large-scale experimentation. In addition, its strong recognition within the open-source community ensures reproducibility and broad relevance. These properties make DeepSeek-R1 an appropriate foundation for controlled analysis of reasoning errors in CoT-style outputs (see Appendix B.2 for comparative results on other models).The following presents the dataset’s construction principles and workflow.
@@ E.1 Evaluation Setting
Model Setting. For the Dataset Construction section, we used the DeepSeek-R1 API [liu2024deepseek] and ChatGPT-4o API [achiam2023gpt] to synthesize data and assist in the manual verification of samples. We evaluated performance based on DeepSeek’s officially released distilled model, DeepSeek-R1-Distill-Qwen-14B [liu2024deepseek] for the Hallucination Detection section. Some detection methods required long-text understanding and processing, such as splitting full CoTs into individual claims, for which we employed DeepSeek-V3 API [liu2024deepseek] and ChatGPT-4o API.
@@ E.2.1 Logit-Based Detection
Logit entropy measures the dispersion of probability mass across the most likely next tokens, making it effective at revealing the model’s uncertainty and exposing potential factual inconsistencies. The model’s output logits are first converted into token‐level perplexity, with higher perplexity indicating greater uncertainty or likelihood of hallucination. For detection, we compute top-k logit entropy by normalizing the entropy over the K most probable tokens, following Sriramanan et al. [sriramanan2024llm], we select an optimal threshold on the validation set, and then apply that threshold during testing to flag hallucinated outputs.
@@ E.2.2 Attention-Based Detection
The attention-kernel score is computed by taking the logarithm of each token’s self-attention weight (i.e., the diagonal entries of the attention matrix) and averaging across tokens. This score effectively measures the model’s degree of “self-focus” during reasoning: consistently high self-attention suggests the model is coherently building on its previous inference, whereas a weakened diagonal focus may indicate hallucination. Following the approach of [sriramanan2024llm], we calculate the attention-kernel score layer by layer and determine, on a validation set, the optimal threshold for each layer. We apply the threshold corresponding to the layer that achieved the best validation performance for testing.
@@ E.2.3 Hidden-State Based Detection
For each sample (question–answer pair), we obtain the hidden representations of each layer in teacher-forcing mode. For each layer’s activation matrix, we compute the centered covariance and perform singular value decomposition (SVD), then average the logarithm of the singular values to derive that layer’s hidden-state score. Higher scores indicate more complete, structured, and reliable internal representations; lower scores often correspond to degraded or incoherent features, potentially signaling hallucination. Following Sriramanan et al. [sriramanan2024llm], we determine the optimal threshold for each layer on a balanced validation set, and during testing, we apply the threshold from the best-performing layer to detect hallucinations in new samples.
@@ E.2.4 HalluciNot (HDM2 model)
The HDM2 model [paudel2025hallucinothallucinationdetectioncontext] detects hallucinations by verifying both context-specific facts and common‐knowledge statements against a fine‐grained taxonomy of LLM outputs. Given a user prompt, an optional context, and an LLM response, it first categorizes each sentence into one of four classes (context‐based, common‐knowledge, enterprise‐specific, or innocuous) using a lightweight classifier. For context‐based claims, it employs a retrieval‐augmented consistency check. For common‐knowledge claims, it probes a frozen backbone LLM’s internal representations to identify contradictions with widely accepted facts. It then produces a document‐level hallucination score and a group of token‐level scores. We took document‐level hallucination scores to evaluate hallucinations.
@@ E.2.5 Claim-Conditioned Perplexity (CCP)
The CCP method quantifies the model’s confidence in factual statements and can be used to detect latent hallucinations in generated text. First, the model’s output is segmented into atomic factual claims, each representing an independent fact unit. We retrieve each token in a claim’s probability distribution over the top-K candidates given the context. Using a natural language inference (NLI) model, we filter these candidates to retain only those that entail or contradict the original token, and compute a normalized ratio of their probabilities as the token’s confidence score. We then aggregate the token-level confidence scores within each claim to obtain an overall uncertainty score for that claim. We average the uncertainty scores across all claims to evaluate an entire text. Finally, we determine an optimal threshold on a validation set to decide whether the text contains hallucinations. We follow the approach described in [fadeeva2024fact] to perform the computations.
@@ E.2.6 SelfCheckGPT
SelfCheckGPT evaluates a model’s self-consistency under diverse random sampling to detect latent hallucinations. Given a user prompt, it first generates a deterministic “main” response at temperature 0, then produces $N$ stochastic samples at temperature 1. The main response is segmented into sentences, and each sentence is compared against all sampled outputs via multiple consistency checks, such as maximal BERTScore similarity inversion, NLI-based contradiction probability, n-gram model token scoring, and prompt-based “Does this hold?” queries. Each check yields a per-sentence uncertainty score, which are aggregated to form both sentence-level and response-level hallucination scores. A threshold tuned on a validation set is then applied to classify sentences or entire responses as hallucinated.We follow the procedure in [manakul2023selfcheckgpt] to perform these calculations.
@@ E.2.7 Semantic Entropy
The Semantic-Entropy score assesses semantic uncertainty by measuring the entropy of the answer distribution obtained through multiple samplings of adversarial prompts. First, the generated text is segmented into atomic claims, and several interrogative prompts are automatically constructed for each claim. Each prompt is then sampled multiple times to gather all answers along with their generation probabilities. Using a bidirectional entailment model, answers are clustered by semantic consistency, and the probabilities within each cluster are summed to form a cluster-level distribution. The information entropy of this distribution is computed as the semantic uncertainty metric for that prompt. We then average the entropies of all prompts associated with the same claim to derive the claim’s overall semantic entropy. Finally, an optimal entropy threshold is determined on the validation set and applied during testing to detect hallucinations. We follow the procedure in [farquhar2024detecting] to perform these calculations.